OpenStudy (anonymous):

3x3y + 12xy - 3x2y - 12y Part A: Rewrite the expression so that the GCF is factored completely. (3 points) Part B: Rewrite the expression completely factored. Show the steps of your work. (4 points) Part C: If the two middle terms were switched so that the expression became 3x3y - 3x2y + 12xy - 12y, would the factored expression no longer be equivalent to your answer in part B? Explain your reasoning. (3 points)

OpenStudy (anonymous):

I'll give you a medal if i get just the answer for a b and c

OpenStudy (anonymous):

@Quinn.Lavnt @Abhisar @anonymous_user @AnswerMyQuestions @Auxuris @animal_lover36 @ashley1211 @ZnappyDooZ @wally15 @Squirrels @Skyz @shereenkhan @swaggd2 @Samus_Aran @sammie2017 @EmogirlAtEmoooocow @edmodo @em0ney @Elsa213

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{\text{is it this }\\~\\ 3x^3y + 12xy - 3x^2y - 12y}\end{align}\)

OpenStudy (anonymous):

eah thats it

OpenStudy (anonymous):

Final result : 3y • (x2 + 4) • (x - 1) Reformatting the input : Changes made to your input should not affect the solution: (1): "x2" was replaced by "x^2". 1 more similar replacement(s).

OpenStudy (anonymous):

all i see are ?????????????? i am lost o.0

OpenStudy (mathmath333):

refresh page

OpenStudy (anonymous):

okay so whats A B and C on that i think 3y(^2+4)(x-1) is the GCF right? idk xD sorry

OpenStudy (anonymous):

i am kind of slow...

OpenStudy (anonymous):

3x3y + 12xy - 3x2y - 12y = 9xy + 12xy - 6xy - 12y = 9xy + 6xy - 12y

OpenStudy (anonymous):

thats part a?

OpenStudy (anonymous):

= 15xy - 12y GCF of 15xy and 12y is 3y

OpenStudy (anonymous):

and we get this 3y(5x - 4)

OpenStudy (anonymous):

that's all :)

OpenStudy (anonymous):

so that all A B and C?

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ A.)\\~\\ 3x^3y + 12xy - 3x^2y - 12y\\~\\ \text{u have to write it such that}\\~\\ \text {their must be a common factor among all}\\\\~\\ =3y(x^3+4x-x^2-4)\\~\\ \text{3y is the common factor here}}\end{align}\)

OpenStudy (anonymous):

ooooh ok ty so much lets move on to B now :)

OpenStudy (anonymous):

u take me wrong @ SouthernRebel101

OpenStudy (anonymous):

i mean to say that's it.

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{\\~\\ B.)~\\~\\ =3y(x^3+4x-x^2-4)\\~\\ \text{for B.) u have to factor it completly}\\~\\ =3y[x(x^2+4)-(x^2+4)]\\~\\ =3y[(x^2+4)(x-1)]}\end{align}\)

OpenStudy (anonymous):

the last one is the final answer?

OpenStudy (anonymous):

for B right?

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{\\~\\ C.)~\\~\\ =3x3y - 3x2y + 12xy - 12y\\~\\ \text{is same as }\\~\\ \color{red}{=3x3y + 12xy- 3x2y - 12y}\\~\\ \text{so exhanging the middle term }\\~\\ \text{won't let any change in the answer }\\~\\ \text{the answer will be the same}}\end{align}\)

OpenStudy (anonymous):

TY